# Controlling the Exchange Field of Surface Spin Impurities via DC Voltages

### v2

@@ Abstract
Comparison of different magnetic impurities– The exchange field model allows to differentiate between an FePc molecule and a Ti atom. Optimizing the simulation parameters to reproduce the the ESR spectra of FePc and Ti measured with the same tip (Fig. 5(a) and (b), respectively) shows good agreement for each species with model parameters summarized in Fig. 5(c). For Ti, we obtained values for $\varepsilon$, $U$, and $\gamma_{\text{T}}$ similar to those in Fig. 4(a) and (b). In contrast, FePc yields significantly different model parameters. When fixing $U_{\text{FePc}}=2.6$ eV based on the $dI/dV$ in Fig. 3(a), the model reproduces the frequency shift with $\varepsilon=-2.032$ eV which matches the negative charge resonance in Fig. 3(a). Allowing all parameters to vary in the fit gives slightly different $U_{\text{FePc}}$ and $\varepsilon$ (see SM, Fig. S9), but both fits indicate that FePc has a much higher Coulomb repulsion and ionization energy than Ti. The energy alignment on the Ag/MgO surface places the unoccupied states closer to the Fermi level, making it easier to access transport at positive bias. This could be a consequence of FePc on Ag/MgO being in an anionic state [13, 45], and has direct implications for the measurement of the ESR signal, as it predicts lower Rabi rates for negative voltages compared to positive ones since $|\varepsilon|\gg\varepsilon+U$  [20, 46]. Finally, the fitted $\gamma_{\text{T}}$ for FePc is substantially larger than for Ti, suggesting a smaller tip-sample distance, consistent with the experiments [33, 10, 16, 45].
Conclusions– In summary, we investigated the influence of DC electric field on the electron spin resonance of individual molecules and metal atoms on a surface. Our model shows that DC bias modulates the exchange field between spin-polarized electrodes and surface spins. This framework, combining exchange and external magnetic fields, explains the observed behaviors including difference between in-plane and out-of-plane magnetic fields, tip angle effects, bistability magnetic tips through polarization inversion, and different responses of two distinct adsorbates. Furthermore, our findings demonstrate the potential of utilizing ESR-STM for characterizing ionization energies of surface spins and revealing fundamental physics of exchange fields stemming from spin-polarized electrodes in surface-spin systems, advancing atomic-scale quantum control through precise DC voltage manipulation.
@@
Acknowledgements– All authors acknowledge support from the Institute for Basic Science under grant IBS-R027-D1. X.Z. acknowledges financial support from the National Natural Science Foundation of China (grant agreement No. 22202074). J.M. received support from National Science Centre of Poland (Grant No. 2020/36/C/ST3/00539). We also thank Nicolás Lorente, Hong Thi Bui, Soo-hyon Phark, Christian Ast, Piotr Kot, Arzhang Ardavan, Piotr Busz, Damian Tomaszewski, Józef Barnaś, and Jürgen König for the fruitful discussions.
@@ I Appendix
FePc molecules, Ti and iron (Fe) atoms were deposited sequentially onto MgO/Ag(100). FePc molecules were deposited at room temperature, while Ti and Fe atoms were deposited onto the cold substrate. The sample was then cooled to a cryostat temperature of 2 K in the STM. The metal atoms and molecules are well-isolated from each other, with a low coverage of approximately 0.05 monolayers. The Ti atoms predominantly adsorb on the oxygen-oxygen bridge site in our work and are referred to simply as Ti. Fe atoms adsorb exclusively on the oxygen site and exhibit a shorter topographic height than Ti atoms. FePc molecules preferentially adsorb with their central Fe atom located on the oxygen site of the MgO lattice. We prepare spin-polarized tips by transferring several Fe atoms onto the tip apex with an applied external magnetic field of $\sim$0.6 T along the out-of-plane direction. All STM and ESR measurements were performed at 2 K with the external magnetic field applied out-of-plane or in-plane.
ESR set-up– The experimental setup for ESR measurements consisted of a signal generator (Keysight, E8257D) providing radio-frequency voltage $V_{\text{RF}}$. DC bias ($V_{\text{DC}}$) was mixed with $V_{\text{RF}}$ through a bias tee and applied to the tip. The sample was grounded. For ESR signal detection, $V_{\text{RF}}$ was modulated at 95 Hz and measured using lock-in amplification. The tunneling current and ESR signal were detected by an electrometer connected to the sample.
@@
To maintain a constant tip-atom distance ($d$) as illustrated in the main text, we adjusted the setpoint current ($I_{\text{set}}$) correspondingly for each $V_{\text{DC}}$, thereby varying the DC bias (more details are provided in SM, Sec. 2). This method allows ESR measurements with active feedback using minimal gain, preventing unexpected thermal drift of the tip height. The ESR spectra are fitted with a Lorentzian to extract the resonance frequency $f_{0}$ for each applied $V_{\text{DC}}$.
@@
Simulations– DFT calculations were performed using plane-wave DFT with pseudopotentials implemented in the Quantum Espresso package [47, 48]. Details of the input parameters are provided in the SM (Sec. 5). Transport simulations were carried out using our own method outlined elsewhere [49, 39, 20]. All solid curves in Figures 3 - 5 are the result of transport simulations with optimized parameters for the impurity, tip angle and electrode-impurity couplings as listed in the respective panels. All simulations were performed at $T=2$ K and we use a fixed ratio of the tip and substrate couplings $\gamma_{\mathrm{S}}=3\gamma_{\mathrm{T}}$ where the unpolarized couplings are $\gamma_{\alpha}=2\pi|t_{\alpha}|^{2}D_{\alpha}$, being $t_{\alpha}$ the tunneling amplitude and $D_{\alpha}$ the spinless density of states per electrode $\alpha$, with $\alpha=\text{T},\text{S}$ representing tip and sample respectively. The density of states is assumed to be constant in the so-called wide-band limit. The spin-polarization of the ferromagnetic tip, defined from the spin-dependence density of states of the tip as $P_{\text{T}}\equiv(D_{T\uparrow}-D_{T\downarrow})/(D_{T\uparrow}+D_{T\downarrow})$, where $D_{T\sigma}=D_{T}\left(1/2+\sigma P_{T}\right)$, was set to $P_{\mathrm{T}}=\pm 0.8$ following [42]. From the model perspective, a lower $P_{\mathrm{T}}$ can be compensated by a higher $\gamma_{\mathrm{T}}$, since both enter Eq. (3) as a product. However, the ESR signals are strongly influenced by the spin-polarization as well as the coupling to tip and substrate. More details of the transport simulations are provided in the SM (Sec. 5).

### v3

@@ Abstract
Comparison of different magnetic impurities– The exchange field model allows to differentiate between an FePc molecule and a Ti atom. Optimizing the simulation parameters to reproduce the ESR spectra of FePc and Ti measured with the same tip (Fig. 4(a) and (b), respectively) shows consistent and reproducible values for each species, see Fig. 4(c). For Ti, we obtain values for $\varepsilon$, $U$, and $\gamma_{\text{T}}$ similar to those in Fig. 3(a) and (b). In contrast, FePc yields significantly different model parameters. When fixing $U=2.6$ eV, based on the $dI/dV$ in Fig. 2(b), the model reproduces the frequency shift with $\varepsilon=-2.032$ eV which matches the negative charge resonance in Fig. 2(b)  [^7] The energy alignment on the Ag/MgO surface places the unoccupied states closer to the Fermi level, making it easier to access via transport at positive bias. This could be a consequence of FePc on Ag/MgO being in an anionic state [28, 34].. Allowing all parameters to vary in the fit gives slightly different $U$ and $\varepsilon$ from those in Fig. 4(c) for FePc [78], but both fits indicate that FePc has a much higher Coulomb repulsion and ionization energy than Ti. Furthermore, while the lack of a strongly non-linear region in our frequency shift data makes a precise determination of $\varepsilon$ and $U$ difficult, our extracted values are in excellent agreement with more recent measurements on the same system which were performed with a $eV_{\mathrm{DC}}>\varepsilon+U$ [40], thereby directly visualizing the non-linear region and confirming the validity of our model and the robustness of our parameter extraction. Importantly, the fitted $\gamma_{\text{T}}$ for FePc is substantially larger than that for Ti, suggesting a smaller tip-sample distance, consistent with experiments [15, 44, 19, 34].
[…]
Figure 5: Unified rescaled behavior of the resonance frequency shift of Ti described by the exchange-field model. (a) Datasets from Fig. 3, Fig. 4, and additional ones (see Supplementary Material [80]) are combined and rescaled to obtain a unified fit using a single set of adjustable parameters: $F=\gamma_{\mathrm{T}}P_{\mathrm{T}}\cos\theta$, $U$, and $\varepsilon$. The dataset acquired with a bistable tip of positive polarization is inverted. The grey shaded area indicates the overall uncertainty of the united fit. (b) Scaling factors $r$ used to align the datasets in (a) and the corresponding fit parameters with uncertainties indicated by the shaded area.
Conclusions– In summary, we investigated how the DC bias influences the electron spin resonance of individual molecules and metal atoms on a surface. Our transport model reveals that the DC bias modulates the exchange field between a spin-polarized electrode and surface spins. This framework, combining exchange and external magnetic fields, captures the observed phenomena, including the difference between in-plane and out-of-plane magnetic fields, tip-angle effects, and the behavior of bistable magnetic tips through polarization inversion. Moreover, the frequency response to the DC bias enables the unambiguous identification of the two adsorbates observed in our experiments. This is supported by extensive statistical analysis across multiple STM tips, highlighting the potential of the exchange field as a tool for highly localized impurity spectroscopy. Altogether, our findings establish the DC bias as an effective parameter for tuning the exchange field in STM-based ESR experiments, overcoming potential limitations of the applicability of the ESR-STM technique. The presented model offers a practical framework for interpreting bias-dependent spin dynamics in nanoscale junctions and can be readily applied to related transport experiments.
@@
Acknowledgments– All authors acknowledge support from the Institute for Basic Science under grant IBS-R027-D1. X.Z. acknowledges financial support from the National Natural Science Foundation of China (grant agreement No. 22202074, 22572060). J.M. received support from National Science Centre of Poland (Grant No. 2020/36/C/ST3/00539). J.R-G acknowledges support from the German Research Foundation [Deutsche Forschungsgemeinschaft (DFG)] under Projects No. 450396347. We also thank Nicolás Lorente, Hong Thi Bui, Soo-hyon Phark, Christian Ast, Piotr Kot, Arzhang Ardavan, Piotr Busz, Damian Tomaszewski, Józef Barnaś, and Jürgen König for engaging discussions.
@@ I Appendix
FePc molecules, Ti and iron (Fe) atoms were deposited sequentially onto MgO/Ag(100). FePc molecules were deposited at room temperature, while Ti and Fe atoms were deposited onto the cold substrate. The sample was then cooled to a cryostat temperature of 1.8 K in the STM. The metal atoms and molecules are well-isolated from each other, with a low coverage of approximately 0.05 monolayers. The Ti atoms predominantly adsorb on the oxygen-oxygen bridge site in our work and are referred to simply as Ti. Fe atoms adsorb exclusively on the oxygen site and exhibit a shorter topographic height than Ti atoms. FePc molecules preferentially adsorb with their central Fe atom located on the oxygen site of the MgO lattice. All measurements were performed on Ti atoms and FePc molecules specifically chosen in defect-free regions of the surface. We prepare spin-polarized tips by transferring several Fe atoms onto the tip apex with an applied external magnetic field of $\sim$0.6 T along the out-of-plane direction. All STM and ESR measurements were performed at 1.8 K with the external magnetic field applied out-of-plane or in-plane.
@@
ESR set-up– The experimental setup for ESR measurements consisted of a signal generator (Keysight, E8257D) providing radio-frequency voltage $V_{\text{RF}}$. The DC bias ($V_{\text{DC}}$) was mixed with $V_{\text{RF}}$ through a bias tee and applied to the tip. The sample was grounded. For ESR signal detection, $V_{\text{RF}}$ was modulated at 95 Hz and measured using lock-in amplification. The tunneling current and ESR signal were detected by an electrometer connected to the sample. To maintain a constant tip-atom distance ($d$) as illustrated in the main text, we adjusted the setpoint current ($I_{\text{set}}$) correspondingly for each $V_{\text{DC}}$, thereby varying the DC bias (more details are provided in [81]). This method allows ESR measurements with active feedback using minimal gain, preventing unexpected thermal drift of the tip height. The ESR spectra are fitted with a Lorentzian to extract the resonance frequency $f_{0}$ for each applied $V_{\text{DC}}$.
@@
Simulations– DFT calculations were performed using plane-wave DFT with pseudopotentials implemented in the Quantum Espresso package [82, 83]. Details of the input parameters are provided in [53]. Transport simulations were carried out using our own method outlined elsewhere [84, 71, 45]. All solid curves in Figures 3-5 are the result of transport simulations with optimized parameters for the impurity, tip angle and electrode-impurity couplings as listed in the respective panels. All simulations were performed at $T=1.8$ K and we use a fixed ratio of the tip and substrate couplings $\gamma_{\mathrm{S}}=3\gamma_{\mathrm{T}}$ where the unpolarized couplings are $\gamma_{\alpha}=2\pi|t_{\alpha}|^{2}D_{\alpha}$, being $t_{\alpha}$ the tunneling amplitude and $D_{\alpha}$ the spinless density of states per electrode $\alpha$, with $\alpha=\text{T},\text{S}$ representing tip and sample, respectively. The density of states is assumed to be constant in the so-called wide-band limit. The spin-polarization of the ferromagnetic tip, defined from the spin-dependence density of states of the tip as $P_{\text{T}}\equiv(D_{T\uparrow}-D_{T\downarrow})/(D_{T\uparrow}+D_{T\downarrow})$, where $D_{T\sigma}=D_{T}\left(1/2+\sigma P_{T}\right)$, was set to $P_{\mathrm{T}}=\pm 0.8$ following [74]. From the model perspective, a lower $P_{\mathrm{T}}$ can be compensated by a higher $\gamma_{\mathrm{T}}$, since both enter Eq. (3) as a product. However, the ESR signals are strongly influenced by the spin-polarization as well as the coupling to tip and substrate. More details of the transport simulations are provided in the Supplemental Materials [68, 64, 78, 77].